Two objects are both negatively charged with 0.02 each and are 70 cm apart. What kind of force exist between them and how much?
1
Expert's answer
2018-11-07T11:08:09-0500

F=1/(4πε_0 ) |q||q|/R^2 →1/(4π∙8,85×〖10〗^(-12) ) 〖0,02〗^2/〖0,7〗^2 =1/(1,11×〖10〗^(-10) ) (4×〖10〗^(-4))/(4,9×〖10〗^(-1) )=7,3×〖10〗^6 N